A firm with annual sales of $8,300,000 increases its inventory turnover from 2.5 to 4.5. How much would the company save annually in interest expense if the cost of carrying the inventory is 7 percent? Round your answer to the nearest dollar.
Answer #1

Current average inventory = 8,300,000/2.5 = 3,320,000

Proposed average inventory = 8,300,000 / 4.5 = 1,844,444.44

Change in inventory = 3,320,000 - 1,844,444.44 = 1,475,555.56

Savings in interest = 1,475,555.56 *7% = 103,289 ( rounded off)